OpenStudy (milrawr):

The density of mercury is 13.6g/mL. How many quarts are in 100 lbs of mercury?

OpenStudy (jiteshmeghwal9):

Volume=mass/density 100lb=454 gram Volume=454/13.6=33.38 ml

OpenStudy (agent0smith):

*1 lb=454 gram

OpenStudy (jiteshmeghwal9):

1 quart implies 940 ml 1 ml implies 1/940 quarts. 33.38 ml implies how much quarts

OpenStudy (milrawr):

@jiteshmeghawal9 wait what?

OpenStudy (milrawr):

that's not what the question is looking for

OpenStudy (agent0smith):

His first post gives you the volume of 1 lb, in mL. You have 100 lb... Once you've done that, then convert mL to quarts using the conversion factor in his second post.

OpenStudy (milrawr):

so... is it 3.5?

OpenStudy (agent0smith):

I didn't work it out. Show your work.

OpenStudy (milrawr):

33.38*100=3338 3338 milliters to quart is 3.573

OpenStudy (milrawr):

i honestly don't know how to figure this out and would really appreciate like an explanation...

OpenStudy (agent0smith):

3.573 quarts looks about right. I assume you know the formula for density = mass/volume, that was used in the first post. You have mass of 100 pounds, which you need to convert to grams, since you have density in g/mL. You plug in the mass and density, and solve for volume. Then you convert volume in mL to quarts.

OpenStudy (mrnood):

Or drop the stupid quarts pounds slugs units for ever and join the rest of the world using the consistent set odf metric SI units

OpenStudy (agent0smith):

^those were my thoughts upon seeing this question with grams, millilitres, quarts, and pounds. What a clusterf---.

OpenStudy (mrnood):

"The density of mercury is 13.6 kg/l" How many litres in 50 kg? 50kg/ 13.6 kg/l =3.68 l Simples

OpenStudy (mrnood):

hmm spose it should strictly be m^3 as volume unit....

OpenStudy (agent0smith):

That's too easy, though. It makes too much sense. There aren't any units with nonsense conversions. Imperial units are to blame for about a 400 million dollar failed landing on Mars, too.